Which of the following matrices is not diagonalizable? OA. [22] 2 0 OB. [22] 22 OC. ſo 21 22] 02 OD. ſo 20] E.
Answer #1

[ ] =L* 21 T2 2 Given that A= 02 [2-2 2 The characteristic matrix is A-AI = 0 2-2 12-2 2 The characteristic equation is|A- A1